### v2

# JuliaQCD: Portable lattice QCD package in Julia language

Abstract
We present JuliaQCD, a new lattice gauge theory codebase developed in the Julia language. Julia is well suited to integrating machine-learning techniques and enables rapid prototyping and execution of algorithms for four-dimensional QCD and other non-Abelian gauge theories. The code leverages LLVM for high-performance execution and supports MPI for parallel computing. Julia’s multiple dispatch provides a flexible and intuitive framework for development. The code implements existing algorithms such as Hybrid Monte Carlo (HMC) and supports many colors and flavors, lattice fermions, smearing techniques, and full QCD simulations. It is designed to run efficiently across various platforms, from laptops to supercomputers, allowing seamless scalability. The codebase is currently available on GitHub at https://github.com/JuliaQCD.
I Introduction
Quantum Chromodynamics (QCD) is the fundamental theory describing the interactions of quarks and gluons, the elementary particles that make up the subatomic world. Because they can be created and annihilated from the vacuum by relativistic processes, they must be treated as a quantum many-body system with an indeterminate particle number. QCD is a quantum field theory that accurately describes our universe.
However, perturbation theory fails in QCD due to the strong coupling. Lattice QCD addresses this challenge by employing lattice regularization, introducing a discrete spacetime cutoff to regulate the path integral. This approach enables the computation of quantum expectation values of physical observables.
Lattice QCD was first formulated by K. G. Wilson in 1974 Wilson (1974), and numerical studies of QCD using this method were initiated by M. Creutz Creutz (1980). This framework enables detailed studies of the QCD vacuum structure and provides precise insights into the Standard Model. Additionally, it facilitates investigations into phenomena such as the quark-gluon plasma and dark-matter candidates.
To formulate lattice QCD, a spacetime lattice cutoff is introduced. This cutoff is not merely a numerical approximation; it fundamentally defines QCD itself. Traditionally, quantum field theory is formulated on a continuous spacetime, but in the presence of interactions such theories suffer from divergences. Although these divergences can be renormalized, they are inherent in the continuum formulation.
In QCD, we aim to calculate the quantum expectation value of a gauge-invariant operator $O$  [^1] For certain purposes, we sometimes calculate the expectation value of a gauge-variant operator. In such cases, gauge fixing is required by Elitzur’s theorem Elitzur (1975).. In the path integral formalism, this is expressed as
$\langle O\rangle=\frac{1}{Z}\int\mathcal{D}U\mathcal{D}\bar{q}\mathcal{D}q\;\mathrm{e}^{-S_{\text{g}}[U]-S_{\text{f}}[U,q,\bar{q}]}O(U,q,\bar{q}),$ (1)
where $q$ and $\bar{q}$ are the quark and antiquark fields, respectively, and $U$ is the gauge field on the lattice. The term $S_{\text{g}}[U]$ represents the gauge field action, and $S_{\text{f}}[U,q,\bar{q}]$ denotes the action for the quarks. The factor $Z$ is a normalization constant ensuring that $\langle 1\rangle=1$. One can obtain observables in QCD by evaluating this integral.
The standard algorithm used for simulating lattice QCD is the Hybrid Monte Carlo (HMC) algorithm Duane et al. (1987). HMC has become the de facto standard because of its efficiency and effectiveness in dealing with the high-dimensional integrals for lattice gauge fields with dynamical fermions. It has been successfully used in a wide range of studies, from calculating the hadron spectrum Ali Khan et al. (2002) and form factors to determining the critical temperature ($T_{c}$) of QCD. The success of lattice QCD in these areas highlights the power of numerical methods in exploring nonperturbative aspects of quantum field theory.
Recent developments in machine learning (ML) have opened new avenues for enhancing lattice QCD calculations. Many researchers are developing new algorithms that integrate ML techniques into lattice QCD simulations. Prototyping is crucial in this context, as we aim to create efficient algorithms for lattice QCD that leverage ML techniques. One of the central trade-offs in this process is between code simplicity and execution speed. While ease of implementation is important, code that is too slow is unusable in practice. Therefore, we focus on optimizing the total time required for lattice QCD studies, which includes both the time spent developing the algorithm and the time required for its execution.
Portability and parallelizability are also crucial considerations. The code must run efficiently on a wide range of platforms, from laptops to supercomputers, and should be optimized to reduce the overall computational time. This allows researchers to take advantage of the available computational resources and scale their simulations as needed.
The Julia language is a new open-source scientific programming language built on LLVM Bezanson et al. (2012). Unlike interpreters, Julia is a compiled language with a just-in-time (JIT) compiler. This allows it to achieve performance comparable to Fortran and C jul , while maintaining the simplicity of Python. Julia’s flexibility and speed make it an ideal choice for developing new lattice QCD algorithms, as it allows for rapid prototyping without sacrificing performance. Thanks to these advantages, Julia is an attractive language for high-energy physics Eschle et al. (2023).
In this paper we introduce a new lattice QCD code, JuliaQCD, written in the Julia language. The motivation behind the development of this codebase is to minimize the combined time of algorithm implementation and execution. Existing languages such as Python are too slow for large-scale lattice QCD simulations, while C++/C/Fortran present challenges in terms of ease of use and development time. Julia strikes the right balance, offering both the speed required for intensive computations and the simplicity needed for fast prototyping. Our goal is to create a flexible and high-performance framework that can be easily adapted to new developments in both lattice QCD and machine learning.
I.1 JuliaQCD project
JuliaQCD is the first open-source lattice QCD code suite in the Julia programming language. (Ultimate) high performance is not within our scope, but we focus on prototyping with good scalability. We developed this code with the following goals:
  1. 1.
    Excellent portability
  2. 2.
    Quick and easy setup
  3. 3.
    Comprehensive suite
  4. 4.
    Highly modifiable
  5. 5.
    Competitive performance
The system offers excellent portability, as it runs smoothly on any machine with Julia installed, without the need for complicated setup processes. It is also quick and easy to set up, enabling users to start working in less than 5 minutes without requiring complex compilation steps. In addition, the system provides a comprehensive suite that supports both full QCD and quenched configuration generation, along with a variety of measurement tools. Its highly modifiable structure makes it ideal for rapid prototyping and experimenting with different approaches. Finally, the system delivers competitive performance, comparable to traditional Fortran 90 codes, ensuring efficiency without sacrificing flexibility.
I.2 Related Work
In the broader landscape, traditional codes such as the MILC code and the Columbia Physics System (CPS) have also played significant roles. Despite their long history, these codes remain functional and are used in various research contexts. MILC, in particular, has been a cornerstone for many researchers, supporting a wide range of computations, including those utilizing GPU architectures mil . CPS continues to be a relevant reference point for its C-based implementation.
FermiQCD Di Pierro (2002) has also contributed to the field. FermiQCD, a collection of C++ classes and parallel algorithms, facilitates fast development of lattice applications and includes optimizations for cluster computing, such as SSE2 instructions. Its focus on distributed processing has allowed it to remain an important tool in parallel lattice QCD.
OpenQCD has established itself as a flexible and advanced package for lattice QCD simulations, particularly in high-performance environments. It supports a wide range of theories, including O(a)-improved Wilson quarks, and includes specialized algorithms such as HMC and SMD. OpenQCD’s flexibility allows for large-scale simulations and efficient use of contemporary processors, making it a widely used tool for complex lattice QCD computations.
The development of lattice QCD codes has been significantly influenced by contributions from various international communities, including notable efforts from Japan. Two key projects, the Lattice Tool Kit in Fortran 90 (LTK) Choe et al. (2002) and Bridge++ Akahoshi et al. (2022), originate from this community. LTK, primarily written in Fortran, has served as a foundational tool, though it is not optimized for contemporary high-performance computing (HPC) environments. In contrast, Bridge++ was developed with a focus on providing a comprehensive and efficient toolset for lattice QCD simulations on modern supercomputers. Both projects reflect Japan’s strong engagement in advancing computational techniques for lattice QCD.
In recent years, there has been a trend toward utilizing C++, as in Bridge++, for its advanced features and improved performance in complex simulations. Grid Boyle et al. (2015) is a prime example of this trend, offering robust support for both CPU and GPU computations. The Grid Python Toolkit (GPT) extends this functionality to Python, allowing for easy integration with data analysis and machine learning workflows. Another significant development is SIMULATeQCD Mazur et al. (2024), which specializes in multi-GPU implementations, enabling high-performance lattice QCD computations. Additionally, QUDA (QCD on CUDA) Clark et al. (2010) serves as a sub-library that provides specialized routines for running lattice QCD computations on NVIDIA GPUs. By leveraging CUDA architectures, QUDA offers substantial speedups for specific operations within larger QCD packages, rather than functioning as a standalone QCD software suite.
The increasing adoption of Python in scientific computing has led to the development of tools such as Lyncs-API, pyQCD Spraggs (2014), and the Grid Python Toolkit (GPT). These projects provide high-level interfaces for lattice QCD calculations, combining Python’s user-friendly syntax with the computational efficiency of C++ backends. Lyncs-API, as discussed in Bacchio et al. (2022), offers a flexible API for a variety of QCD computations, while pyQCD similarly integrates Python and C++ for efficient lattice QCD research. This landscape illustrates the evolution of lattice QCD software from traditional Fortran and C implementations to modern, multi-language frameworks designed for contemporary HPC environments. LatticeQCD.jl contributes to this ongoing evolution by leveraging Julia’s capabilities, providing a new, efficient, and accessible option for researchers.
II Software description
We support lattice gauge theory in 4-dimensional Euclidean spacetime. We cover most standard techniques. We also implement a wizard for parameter files. Our code, LatticeQCD.jl, has the following functionalities.
Gauge fields
  • Optimized $\mathrm{SU}(2)$, $\mathrm{SU}(3)$
  • General $\mathrm{SU}(N_{c})$
Fermion actions
  • Wilson (2 flavors)
  • Staggered fermions (1-8 tastes)
  • Standard domain-wall (2 flavors)
Measurements
  • Plaquette
  • Polyakov loop
  • Chiral condensates (Wilson, staggered)
  • Momentum-projected pion correlator (Wilson, staggered)
  • $R\times T$ Wilson loop
  • Energy density
  • Topological charge (plaquette, clover, and $\mathcal{O}(a^{2})$-improved definition)
  • Load & measurement mode (load and measure all configurations in a directory)
Smearing
  • Stout
  • Gradient flow for a generic action
Configuration generation algorithms
  • Cold/Hot start for $\mathrm{SU}(N_{c})$. One instanton configuration for $\mathrm{SU}(2)$
  • Heatbath for $\mathrm{SU}(N_{c})$ & overrelaxation for a general gauge action
  • Even-odd heatbath for the plaquette action
  • Quenched HMC with $\mathrm{SU}(N_{c})$ for a general gauge action
  • HMC (2-flavor Wilson) for $\mathrm{SU}(N_{c})$ with a general gauge action
  • HMC (4-taste staggered fermions) for $\mathrm{SU}(N_{c})$ with a general gauge action
  • RHMC (any-flavor staggered) for $\mathrm{SU}(N_{c})$ with a general gauge action
  • Stout-smeared dynamical fermions for $\mathrm{SU}(N_{c})$
  • Self-learning HMC with the plaquette action
I/O for gauge configurations
  • ILDG format (binary)
  • JLD format (Julia’s default binary file, HDF5-based)
  • Bridge++ text file (Bridgetext)
If one specifies other than $N_{f}=4,8$ with the staggered-fermion HMC, RHMC is automatically used. For a machine with Apple Silicon, $N_{f}=1-8$ is available. Parallelization is supported by MPI, which can be used with Gaugefields.jl and LatticeDiracOperators.jl (see Appendix B.4). We perform consistency checks against several references. See Appendix E.
II.1 Code Structure
JuliaQCD is the project name, and it is composed of several packages.
  • LatticeQCD.jl : LatticeQCD.jl is a wrapper of the following packages. The interface of this package is user-friendly.
  • Wilsonloop.jl: Wilsonloop.jl helps handle Wilson loops and generic Wilson lines in any $N_{c}$ and dimensions. Wilson lines can be defined symbolically.
  • Gaugefields.jl: Gaugefields.jl is a package for lattice $\mathrm{SU}(N_{c})$ gauge fields. It handles gauge fields (links), gauge actions with MPI, and autograd. It can generate quenched configurations.
  • LatticeDiracOperators.jl: LatticeDiracOperators.jl is a package for Dirac operators and fermions on the lattice. It handles pseudofermion fields with various lattice Dirac operators, and fermion actions with MPI. It can generate configurations with dynamical fermions.
  • QCDMeasurements.jl: QCDMeasurements.jl is a package for measuring physical quantities. It includes measurements for basic quantities like chiral condensates, plaquettes, pion correlators, and topological charge with several definitions. It also includes the gradient flow with several actions.
Figure 1: Structure of the packages (JuliaQCD).
III Usage of LatticeQCD.jl
We present example code to perform lattice QCD calculations using LatticeQCD.jl. One can start a lattice QCD calculation in 5 steps.
  1. 1.
    Install the Julia language from JuliaLang.org.
  2. 2.
    In Julia REPL, press the ] key to enter the package mode and type:
    1add LatticeQCD
    then press the “return” key. Press the “backspace” key (or “delete” key on Mac) to exit package mode. The latest version is available via add LatticeQCD#master. All dependencies will be resolved automatically.
  3. 3.
    Include the package with:
    1using LatticeQCD
  4. 4.
    Make a parameter file with the wizard:
    1run_wizard()
    Choose parameters.
  5. 5.
    Start the simulation with the created parameter file:
    1run_LQCD("my_parameters.toml")
    One will obtain results. One can write a parameter file.
III.1 User interfaces
We support the following two user interfaces for LatticeQCD.jl:
  1. 1.
    Julia REPL interface (for beginners, just after the lattice QCD textbook)
  2. 2.
    General interface (experience with another code, for batch jobs, customized purposes)
Usage 1 is explained above.
For Usage 2, in Julia REPL, press the ] key to enter the package mode and type:
1add LatticeQCD
Then, LatticeQCD.jl will be installed. The “PARAMETER_FILE” can be created through the wizard. To use the wizard in the shell, please write the following code and save it as wizard.jl:
1using LatticeQCD
2run_wizard()
Then, one can run the wizard:
1julia wizard.jl
Please write the following code and save it as run.jl:
1using LatticeQCD
2run_LQCD(ARGS[1])
Then, one can execute it like this:
1julia run.jl PARAMETER_FILE
IV Conclusion and Outlook
In this work, we introduce JuliaQCD, a new and flexible lattice QCD codebase developed in the Julia programming language. By leveraging Julia’s high-performance just-in-time compilation and ease of use, this project aims to strike a balance between rapid prototyping and computational efficiency. JuliaQCD offers a range of functionalities, including support for full QCD simulations and various smearing techniques, while maintaining scalability across different computational environments from personal systems to high-performance computing clusters.
Our approach is motivated by the need to reduce the time required for both algorithm implementation and execution, allowing researchers to experiment with and refine techniques without sacrificing performance. The portability and ease of setup provided by JuliaQCD make it an ideal tool for researchers working at the intersection of lattice QCD and machine learning, where quick adaptation to new algorithms and methodologies is critical.
While the performance of JuliaQCD is comparable to established Fortran codes, its focus remains on flexibility and usability, providing a framework that can be easily extended to incorporate modern computational techniques. Future work will explore further optimizations, including deeper integration with GPU architectures and advanced machine learning techniques to further enhance the capabilities of lattice QCD simulations.
JuliaQCD represents an important step in the evolution of lattice QCD software, contributing to the broader trend of adopting more modern, user-friendly languages without compromising on performance. We anticipate that this work will help advance both the development and application of lattice QCD, facilitating new discoveries in Quantum Chromodynamics and related fields.
Acknowledgements.
The authors thank to the Lattice Tool Kit written in Fortran 90. The authors thank to Taku Izubuchi, Issaku Kanamori, Okuto Morikawa, Satoshi Terasaki and Hiromasa Watanabe for comments and contributions. The work of authors was partially supported by JSPS KAKENHI Grant Numbers 20K14479, 22K03539, 22H05112, and 22H05111, and MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Simulation for basic science: approaching the new quantum era; Grant Number JPMXP1020230411, and Search for physics beyond the standard model using large-scale lattice QCD simulation and development of AI technology toward next-generation lattice QCD; Grant Number JPMXP1020230409).
Appendix A Wilsonloop.jl
A gauge action is constructed by gauge invariant objects, Wilson loops, in discretized spacetime. Wilsonloop.jl helps us to treat with the Wilson loops and generic Wilson lines in any Nc and dimensions. Wilsonloop.jl has the following functionalities:
  • From a symbolic definition of Wilson lines, this returns $\mathrm{SU}(N_{c})$-valued Wilson lines as objects
  • Constructing all staples from given symbolic Wilson lines
  • Constructing derivatives of given symbolic Wilson lines (auto-grad for $\mathrm{SU}(N_{c})$ variables)
We can easily generate a plaquette shown as
1println("plaq")
2plaq = make_plaq()
3display(plaq)
The staple of the plaquette is given as
1for $\mu$=1:4
2 println($\mu$ = $\mu$")
3 staples = make_plaq_staple($\mu$)
4 display(staples)
5end
An arbitrary Wilson loop is constructed as
1loop = [(1,+1),(2,+1),(1,-1),(2,-1)]
2println(loop)
3w = Wilsonline(loop)
4println("P: ")
5show(w)
Its adjoint is calculated as
1println("P^+: ")
2show(w’)
The derivative of the lines $dw/dU_{\mu}$ is calculated as
1for $\mu$=1:4
[…]
3 for i=1:length(dU)
4 show(dU[i])
5 end
6end
Note that the derivative is a rank-4 tensor. The derivatives are usually used for making the smearing of the gauge fields (Stout smearing can be used in Gaugefields.jl).
Appendix B Gaugefields.jl
This package has following functionalities
  • $\mathrm{SU}(N_{c})$ (Nc ¿ 1) gauge fields in 2 or 4 dimensions with arbitrary actions.
  • $Z(N_{c})$ 2-form gauge fields in 4 dimensions, which are given as ’t Hooft flux.
  • U(1) gauge fields in 2 dimensions with arbitrary actions.
  • Configuration generation
    • Heatbath
    • quenched Hybrid Monte Carlo
    • quenched Hybrid Monte Carlo being subject to ’t Hooft twisted b.c. with external (non-dynamical) $Z(N_{c})$ 2-form gauge fields  [^2] Thanks for O. Morikawa.
    • quenched Hybrid Monte Carlo for $\mathrm{SU}(N_{c})$$Z(N_{c})$ gauge theory with dynamical $Z(N_{c})$ 2-form gauge fields
  • Gradient flow via RK3
    • Yang-Mills gradient flow
    • Yang-Mills gradient flow being subject to ’t Hooft twisted b.c.
    • Gradient flow for $\mathrm{SU}(N_{c})$$Z(N_{c})$ gauge theory
  • I/O: ILDG and Bridge++ formats are supported (c-lime will be installed implicitly with CLIME_jll )
  • MPI parallel computation(experimental)
    • quenched HMC with MPI being subject to ’t Hooft twisted b.c.
In addition, this supports followings
  • Autograd for functions with $\mathrm{SU}(N_{c})$ variables
  • Stout smearing
  • Stout force via backpropagation
We note that Autograd can be worked for general Wilson lines except for ones have overlaps.
B.1 File loading and saving
B.1.1 ILDG format
We can use ILDG format, one of standard formats for configurations. We can read ILDG format as follows:
1using Gaugefields
2
3NX = 4
4NY = 4
5NZ = 4
6NT = 4
7NC = 3
8Nwing = 1
9Dim = 4
10U = Initialize_Gaugefields(NC,Nwing,NX,NY,NZ,NT,condition = "cold")
11filename = "hoge.ildg"
12ildg = ILDG(filename)
13i = 1
14L = [NX,NY,NZ,NT]
15load_gaugefield!(U,i,ildg,L,NC)
With the use of the configuration, we can calculate the plaquette:
1temp1 = similar(U[1])
2temp2 = similar(U[1])
3
4comb = 6
5factor = 1/(comb*U[1].NV*U[1].NC)
6@time plaq_t = calculate_Plaquette(U,temp1,temp2)*factor
7println("plaq_t = \$plaq_t")
8poly = calculate_Polyakov_loop(U,temp1,temp2)
9println("polyakov loop = \$(real(poly)) \$(imag(poly))")
We can write a configuration as the ILDG format as follows:
1filename = "hoge.ildg"
2save_binarydata(U,filename)
B.1.2 Text format for Bridge++
Gaugefields.jl also supports a text format for Bridge++. A file loading and saving are expressed as follows:
1using Gaugefields
2filename = "testconf.txt"
3load_BridgeText!(filename,U,L,NC)
1filename = "testconf.txt"
2save_textdata(U,filename)
B.2 Gradient flow
To smear Gauge fields is important in LatticeQCD. We show the codes of the Lüscher’s gradient flow as follows.
1NX = 4
2NY = 4
3NZ = 4
4NT = 4
5Nwing = 0
6NC = 3
7
8U = Initialize_Gaugefields(NC, Nwing, NX, NY, NZ, NT, condition="hot")
9
10temp1 = similar(U[1])
11temp2 = similar(U[1])
12comb = 6
13factor = 1 / (comb * U[1].NV * U[1].NC)
14
15g = Gradientflow(U)
16for itrj = 1:100
17 flow!(U, g)
18 @time plaq_t = calculate_Plaquette(U, temp1, temp2) * factor
19 println("\$itrj plaq_t = \$plaq_t")
20 poly = calculate_Polyakov_loop(U, temp1, temp2)
21 println("\$itrj polyakov loop = \$(real(poly)) \$(imag(poly))")
22end
B.3 Hybrid Monte Carlo
With the use of the Gaugefields.jl, we can easily make the Hybrid Monte Carlo method as following.
1using Random
2using Gaugefields
3using LinearAlgebra
4
5function calc_action(gauge_action, U, p)
6 NC = U[1].NC
7 Sg = -evaluate_GaugeAction(gauge_action, U) / NC
8 Sp = p * p / 2
9 S = Sp + Sg
10 return real(S)
11end
12
13function MDstep!(gauge_action, U, p, MDsteps, Dim, Uold, temp1, temp2)
14 $\Delta$$\tau$ = 1.0 / MDsteps
15 gauss_distribution!(p)
16 Sold = calc_action(gauge_action, U, p)
17 substitute_U!(Uold, U)
18
19 for itrj = 1:MDsteps
20 U_update!(U, p, 0.5, $\Delta$$\tau$, Dim, gauge_action)
21
22 P_update!(U, p, 1.0, $\Delta$$\tau$, Dim, gauge_action, temp1, temp2)
23
24 U_update!(U, p, 0.5, $\Delta$$\tau$, Dim, gauge_action)
25 end
26 Snew = calc_action(gauge_action, U, p)
27 println("Sold = \$Sold, Snew = \$Snew")
28 println("Snew - Sold = \$(Snew-Sold)")
29 ratio = min(1, exp(-Snew + Sold))
30 if rand() > ratio
31 substitute_U!(U, Uold)
32 return false
33 else
34 return true
35 end
36end
37
38function U_update!(U, p, $\epsilon$, $\Delta$$\tau$, Dim, gauge_action)
39 temps = get_temporary_gaugefields(gauge_action)
40 temp1 = temps[1]
41 temp2 = temps[2]
42 expU = temps[3]
43 W = temps[4]
44 for $\mu$ = 1:Dim
45 exptU!(expU, $\epsilon$ * $\Delta$$\tau$, p[$\mu$], [temp1, temp2])
46 mul!(W, expU, U[$\mu$])
47 substitute_U!(U[$\mu$], W)
48 end
49end
50
[…]
52 NC = U[1].NC
53 temp = temp1
[…]
55 factor = -$\epsilon$ * $\Delta$$\tau$ / (NC)
56 for $\mu$ = 1:Dim
[…]
59 Traceless_antihermitian_add!(p[$\mu$], factor, temp)
60 end
61end
62function HMC_test_4D(NX, NY, NZ, NT, NC, $\beta$)
63 Dim = 4
64 Nwing = 0
65 Random.seed!(123)
66 U = Initialize_Gaugefields(NC, Nwing, NX, NY, NZ, NT, condition="hot", randomnumber="Reproducible")
67 println(typeof(U))
68 temp1 = similar(U[1])
69 temp2 = similar(U[1])
70 comb = 6 #4*3/2
71 factor = 1 / (comb * U[1].NV * U[1].NC)
72 @time plaq_t = calculate_Plaquette(U, temp1, temp2) * factor
73 println("0 plaq_t = \$plaq_t")
74 poly = calculate_Polyakov_loop(U, temp1, temp2)
75 println("0 polyakov loop = \$(real(poly)) \$(imag(poly))")
76
77 gauge_action = GaugeAction(U)
78 plaqloop = make_loops_fromname("plaquette")
79 append!(plaqloop, plaqloop’)
80 $\beta$ = $\beta$ / 2
81 push!(gauge_action, $\beta$, plaqloop)
82
83 p = initialize_TA_Gaugefields(U) #This is a traceless-antihermitian gauge fields. This has NC^2-1 real coefficients.
84 Uold = similar(U)
85 substitute_U!(Uold, U)
86 MDsteps = 100
87 numaccepted = 0
88
89 numtrj = 10
90 for itrj = 1:numtrj
91 t = @timed begin
92 accepted = MDstep!(gauge_action, U, p, MDsteps, Dim, Uold, temp1, temp2)
93 end
94 if get_myrank(U) == 0
95 println("elapsed time for MDsteps: \$(t.time) [s]")
96 end
97 numaccepted += ifelse(accepted, 1, 0)
98 if itrj % 10 == 0
99 @time plaq_t = calculate_Plaquette(U, temp1, temp2) * factor
100 println("\$itrj plaq_t = \$plaq_t")
101 poly = calculate_Polyakov_loop(U, temp1, temp2)
102 println("\$itrj polyakov loop = \$(real(poly)) \$(imag(poly))")
103 println("acceptance ratio ", numaccepted / itrj)
104 end
105 end
106 return plaq_t, numaccepted / numtrj
107end
108
109function main()
110 $\beta$ = 5.7
111 NX = 8
112 NY = 8
113 NZ = 8
114 NT = 8
115 NC = 3
116 HMC_test_4D(NX, NY, NZ, NT, NC, $\beta$)
117end
118main()
B.4 HMC with MPI
Here, we show the HMC with MPI. the REPL and Jupyter notebook can not be used when one wants to use MPI. At first, in Julia REPL in the package mode,
1add MPIPreferences
and
1using MPIPreferences
2MPIPreferences.use_system_binary()
With the use of
1add MPI
we can use MPI in Julia. We show the sample code:
1
2using Random
3using Gaugefields
4using LinearAlgebra
5using MPI
6
7if length(ARGS) < 5
8 error("USAGE: ","""
9 mpiexecjl -np 2 exe.jl 1 1 1 2 true
10 """)
11end
12const pes = Tuple(parse.(Int64,ARGS[1:4]))
13const mpi = parse(Bool,ARGS[5])
14
15function calc_action(gauge_action,U,p)
16 NC = U[1].NC
17 Sg = -evaluate_GaugeAction(gauge_action,U)/NC #evaluate_Gauge_action(gauge_action,U) = tr(evaluate_Gaugeaction_untraced(gauge_action,U))
18 Sp = p*p/2
19 S = Sp + Sg
20 return real(S)
21end
22
23function MDstep!(gauge_action,U,p,MDsteps,Dim,Uold,temp1,temp2)
24 $\Delta$$\tau$ = 1.0/MDsteps
25 gauss_distribution!(p)
26 Sold = calc_action(gauge_action,U,p)
27 substitute_U!(Uold,U)
28
29 for itrj=1:MDsteps
[…]
31
[…]
33
[…]
35 end
36 Snew = calc_action(gauge_action,U,p)
37 if get_myrank(U) == 0
38 println("Sold = \$Sold, Snew = \$Snew")
39 println("Snew - Sold = \$(Snew-Sold)")
40 end
41 ratio = min(1,exp(-Snew+Sold))
42 r = rand()
43 if mpi
44 r = MPI.bcast(r, 0, MPI.COMM_WORLD)
45 end
46 #ratio = min(1,exp(Snew-Sold))
47 if r > ratio
48 substitute_U!(U,Uold)
49 return false
50 else
51 return true
52 end
53end
54
[…]
56 temps = get_temporary_gaugefields(gauge_action)
57 temp1 = temps[1]
58 temp2 = temps[2]
59 expU = temps[3]
60 W = temps[4]
61
62 for $\mu$=1:Dim
[…]
64 mul!(W,expU,U[$\mu$])
65 substitute_U!(U[$\mu$],W)
66
67 end
68end
69
[…]
71 NC = U[1].NC
72 temp = temp1
[…]
74 factor = -$\epsilon*$$\Delta$$\tau$/(NC)
75
76 for $\mu$=1:Dim
[…]
79 Traceless_antihermitian_add!(p[$\mu$],factor,temp)
80 end
81end
82
83
[…]
85 Dim = 4
86 Nwing = 0
87 Random.seed!(123)
88 if mpi
89 PEs = pes#(1,1,1,2)
90 U = Initialize_Gaugefields(NC,Nwing,NX,NY,NZ,NT,condition = "hot",mpi=true,PEs = PEs,mpiinit = false)
91 else
92 U = Initialize_Gaugefields(NC,Nwing,NX,NY,NZ,NT,condition = "hot")
93 end
94
95 if get_myrank(U) == 0
96 println(typeof(U))
97 end
98
99 temp1 = similar(U[1])
100 temp2 = similar(U[1])
101
102 if Dim == 4
103 comb = 6 #4*3/2
104 elseif Dim == 3
105 comb = 3
106 elseif Dim == 2
107 comb = 1
108 else
109 error("dimension \$Dim is not supported")
110 end
111
112 factor = 1/(comb*U[1].NV*U[1].NC)
113
114 @time plaq_t = calculate_Plaquette(U,temp1,temp2)*factor
115 if get_myrank(U) == 0
116 println("0 plaq_t = \$plaq_t")
117 end
118 poly = calculate_Polyakov_loop(U,temp1,temp2)
119 if get_myrank(U) == 0
120 println("0 polyakov loop = \$(real(poly)) \$(imag(poly))")
121 end
122
123 gauge_action = GaugeAction(U)
124 plaqloop = make_loops_fromname("plaquette")
125 append!(plaqloop,plaqloop’)
126 $\beta$ = $\beta$/2
[…]
128
129 p = initialize_TA_Gaugefields(U) #This is a traceless-antihermitian gauge fields. This has NC^2-1 real coefficients.
130 Uold = similar(U)
131 substitute_U!(Uold,U)
132 MDsteps = 100
133 temp1 = similar(U[1])
134 temp2 = similar(U[1])
135 comb = 6
136 factor = 1/(comb*U[1].NV*U[1].NC)
137 numaccepted = 0
138
139 numtrj = 100
140 for itrj = 1:numtrj
141 t = @timed begin
142 accepted = MDstep!(gauge_action,U,p,MDsteps,Dim,Uold,temp1,temp2)
143 end
144 if get_myrank(U) == 0
145 println("elapsed time for MDsteps: \$(t.time) [s]")
146 end
147 numaccepted += ifelse(accepted,1,0)
148
149 if itrj % 10 == 0
150 plaq_t = calculate_Plaquette(U,temp1,temp2)*factor
151 if get_myrank(U) == 0
152 println("\$itrj plaq_t = \$plaq_t")
153 end
154 poly = calculate_Polyakov_loop(U,temp1,temp2)
155 if get_myrank(U) == 0
156 println("\$itrj polyakov loop = \$(real(poly)) \$(imag(poly))")
157 println("acceptance ratio ",numaccepted/itrj)
158 end
159 end
160 end
161 return plaq_t,numaccepted/numtrj
162
163end
164
165function main()
166 $\beta$ = 5.7
167 NX = 8
168 NY = 8
169 NZ = 8
170 NT = 8
171 NC = 3
[…]
173end
174main()
The command is like:
1mpiexecjl -np 2 julia mpi_sample.jl 1 1 1 2 true
We can also use MPI in LatticeDiracOperators.jl.
Appendix C LatticeDiracOperators.jl
LatticeDiracOperators.jl handles fermions on a lattice. This package have the following functionalities:
  • Constructing actions and its derivative for staggered fermion with 1-8 tastes with the use of the rational HMC
  • Constructing actions and its derivative for Wilson fermion
  • Constructing actions and its derivative for Standard Domainwall fermion
  • Hybrid Monte Carlo method with fermions
With the use of the Gaugefields.jl, we can also do the HMC with stout smearing. This package can be regarded as the additional package of the Gaugefields.jl to treat with lattice fermions.
C.0.1 Definition of pseudo-fermion fields
The pseudo-fermion fields can be defined as
1using Gaugefields
2using LatticeDiracOperators
3
4NX = 4
5NY = 4
6NZ = 4
7NT = 4
8Nwing = 0
9Dim = 4
10NC = 3
11
12U = Initialize_4DGaugefields(NC,Nwing,NX,NY,NZ,NT,condition = "cold")
13x = Initialize_pseudofermion_fields(U[1],"Wilson")
Here, x is a pseudo fermion fields for Wilson Dirac operator. The element of x is x[ic,ix,iy,iz,it,ialpha]. ic is an index of the color. ialpha is the internal degree of the gamma matrix. The staggered fermions can be defined as
1x = Initialize_pseudofermion_fields(U[1],"staggered")
If one wants to obtain the Gaussian distributed pseudo-fermions, the code is written as
1gauss_distribution_fermion!(x)
C.1 Definition of Dirac operators
The Dirac operators are important basic parts in lattice QCD simulations. The Wilson Dirac operator can be defined as
1params = Dict()
2params["Dirac_operator"] = "Wilson"
3params[$\kappa$"] = 0.141139
4params["eps_CG"] = 1.0e-8
5params["verbose_level"] = 2
6D = Dirac_operator(U,x,params)
We can treat the Dirac operator as a matrix. Thus, we can apply the Dirac operator to the pseudo-fermion fields as follows.
1using LinearAlgebra
2y = similar(x)
3mul!(y,D,x)
And we can solve the equation $Dx=b$:
1solve_DinvX!(y,D,x)
The convergence property can be seen by setting ”verbose_level” flag:
1params["verbose_level"] = 3
2D = Dirac_operator(U,x,params)
3gauss_distribution_fermion!(x)
4solve_DinvX!(y,D,x)
5println(y[1,1,1,1,1,1])
The adjoint of the Dirac operator $D^{\dagger}$ and $D^{\dagger}D$ operator can be defined as
1gauss_distribution_fermion!(x)
2solve_DinvX!(y,D’,x)
3println(y[1,1,1,1,1,1])
4DdagD = DdagD_operator(U,x,params)
5gauss_distribution_fermion!(x)
6solve_DinvX!(y,DdagD,x)
7println(y[1,1,1,1,1,1])
We can similarly define the Dirac operator for the staggered fermion as follows:
1x = Initialize_pseudofermion_fields(U[1],"staggered")
2gauss_distribution_fermion!(x)
3params = Dict()
4params["Dirac_operator"] = "staggered"
5params["mass"] = 0.1
6params["eps_CG"] = 1.0e-8
7params["verbose_level"] = 2
8D = Dirac_operator(U,x,params)
9y = similar(x)
10mul!(y,D,x)
11println(y[1,1,1,1,1,1])
12solve_DinvX!(y,D,x)
13println(y[1,1,1,1,1,1])
The ”tastes” of the staggered fermion is defined in the action.
C.2 Definition of fermion actions
With the use of the LatticeDiracOperators.jl, we can define actions for pseudo-fermions. The sample codes are written as
1NX = 4
2NY = 4
3NZ = 4
4NT = 4
5Nwing = 0
6Dim = 4
7NC = 3
8
9U = Initialize_4DGaugefields(NC, Nwing, NX, NY, NZ, NT, condition="cold")
10x = Initialize_pseudofermion_fields(U[1], "Wilson")
11gauss_distribution_fermion!(x)
12
13params = Dict()
14params["Dirac_operator"] = "Wilson"
15params[$\kappa$"] = 0.141139
16params["eps_CG"] = 1.0e-8
17params["verbose_level"] = 2
18
19D = Dirac_operator(U, x, params)
20
21parameters_action = Dict()
22fermi_action = FermiAction(D, parameters_action)
Then, the fermion action with given pseudo-fermion fields is evaluated as
1Sfnew = evaluate_FermiAction(fermi_action,U,x)
2println(Sfnew)
We can also calculate the derivative of the fermion action $dSf/dU$ as
[…]
In the case of the staggered fermion, we can choose ”taste”. The action is defined as
1x = Initialize_pseudofermion_fields(U[1],"staggered")
2gauss_distribution_fermion!(x)
3params = Dict()
4params["Dirac_operator"] = "staggered"
5params["mass"] = 0.1
6params["eps_CG"] = 1.0e-8
7params["verbose_level"] = 2
8D = Dirac_operator(U,x,params)
9
10Nf = 2
11
12println("Nf = \$Nf")
13parameters_action = Dict()
14parameters_action["Nf"] = Nf
15fermi_action = FermiAction(D,parameters_action)
16Sfnew = evaluate_FermiAction(fermi_action,U,x)
17println(Sfnew)
[…]
This package uses the RHMC techniques to consider the tastes.
C.3 Hybrid Monte Carlo
We show a sample code for the Hybrid Monte Carlo method with pseudo-fermion fields. The codes are written as
1using Gaugefields
2using LatticeDiracOperators
3using LinearAlgebra
4using InteractiveUtils
5using Random
6
[…]
8 p = initialize_TA_Gaugefields(U) #This is a traceless-antihermitian gauge fields. This has NC^2-1 real coefficients.
9 Uold = similar(U)
10 substitute_U!(Uold,U)
11 MDsteps = 10
12 temp1 = similar(U[1])
13 temp2 = similar(U[1])
14 comb = 6
15 factor = 1/(comb*U[1].NV*U[1].NC)
16 numaccepted = 0
17 Random.seed!(123)
18
19 numtrj = 10
20 for itrj = 1:numtrj
[…]
22 numaccepted += ifelse(accepted,1,0)
23
24 plaq_t = calculate_Plaquette(U,temp1,temp2)*factor
25 println("\$itrj plaq_t = \$plaq_t")
26 println("acceptance ratio ",numaccepted/itrj)
27 end
28end
29
30function calc_action(gauge_action,U,p)
31 NC = U[1].NC
32 Sg = -evaluate_GaugeAction(gauge_action,U)/NC
33 Sp = p*p/2
34 S = Sp + Sg
35 return real(S)
36end
37
38
[…]
40 $\Delta$$\tau$ = 1/MDsteps
41 NC,_,NN... = size(U[1])
42
43 gauss_distribution!(p)
44
45 substitute_U!(Uold,U)
46 gauss_sampling_in_action!($\xi$,U,fermi_action)
[…]
48 Sfold = real(dot($\xi$$\xi$))
49 println("Sfold = \$Sfold")
50
51 Sold = calc_action(gauge_action,U,p) + Sfold
52 println("Sold = ",Sold)
53
54 for itrj=1:MDsteps
[…]
56
[…]
59
[…]
61 end
[…]
63 println("Sfnew = \$Sfnew")
64 Snew = calc_action(gauge_action,U,p) + Sfnew
65
66 println("Sold = \$Sold, Snew = \$Snew")
67 println("Snew - Sold = \$(Snew-Sold)")
68 accept = exp(Sold - Snew) >= rand()
69 if accept != true #rand() > ratio
70 substitute_U!(U,Uold)
71 return false
72 else
73 return true
74 end
75end
76
[…]
78 temps = get_temporary_gaugefields(gauge_action)
79 temp1 = temps[1]
80 temp2 = temps[2]
81 expU = temps[3]
82 W = temps[4]
83
84 for $\mu$=1:Dim
[…]
86 mul!(W,expU,U[$\mu$])
87 substitute_U!(U[$\mu$],W)
88
89 end
90end
91
[…]
93 NC = U[1].NC
94 temps = get_temporary_gaugefields(gauge_action)
[…]
96 factor = -$\epsilon*$$\Delta$$\tau$/(NC)
97
98 for $\mu$=1:Dim
[…]
101 Traceless_antihermitian_add!(p[$\mu$],factor,temps[1])
102 end
103end
104
[…]
106 temps = get_temporary_gaugefields(gauge_action)
[…]
108 factor = -$\epsilon*$$\Delta$$\tau$
[…]
110
111 for $\mu$=1:Dim
[…]
113 end
114end
115
116function test1()
117 NX = 4
118 NY = 4
119 NZ = 4
120 NT = 4
121 Nwing = 0
122 Dim = 4
123 NC = 3
124 U = Initialize_4DGaugefields(NC,Nwing,NX,NY,NZ,NT,condition = "cold")
125 gauge_action = GaugeAction(U)
126 plaqloop = make_loops_fromname("plaquette")
127 append!(plaqloop,plaqloop’)
128 $\beta$ = 5.5/2
[…]
130 show(gauge_action)
131 x = Initialize_pseudofermion_fields(U[1],"Wilson")
132
133 params = Dict()
134 params["Dirac_operator"] = "Wilson"
135 params[$\kappa$"] = 0.141139
136 params["eps_CG"] = 1.0e-8
137 params["verbose_level"] = 2
138 D = Dirac_operator(U,x,params)
139 parameters_action = Dict()
140 fermi_action = FermiAction(D,parameters_action)
141 y = similar(x)
142 MDtest!(gauge_action,U,Dim,fermi_action,x,y)
143end
144test1()
We can easily switch the Wilson fermion to the staggered fermions.
Appendix D QCDMeasurements.jl
This package has following functionalities
  • Plaquette measurement.
  • Polyakov loop measurement.
  • Pion correlator measurement.
  • Chiral condensate measurement.
  • Topological charge measurement.
  • Energy density measurement.
  • Wilson loop measurement
D.1 Sample
To measure variables, one have to make an instance
1m_plaq = Plaquette_measurement(U)
By using this m_plaq, one can measure and get the plaquette:
1plaq = get_value(measure(m_plaq,U))
The example is shown as follows.
1using QCDMeasurements
2using Gaugefields
3function test()
4 println("SU3test")
5 NX = 4
6 NY = 4
7 NZ = 4
8 NT = 4
9 Nwing = 0
10 Dim = 4
11 NC = 3
12
13 U = Initialize_4DGaugefields(NC,Nwing,NX,NY,NZ,NT,condition = "cold")
14 filename = "testconf.txt"
15 L = [NX,NY,NZ,NT]
16 load_BridgeText!(filename,U,L,NC)
17 m_plaq = Plaquette_measurement(U)
18 m_poly = Polyakov_measurement(U)
19
20 plaq = get_value(measure(m_plaq,U))
21 poly = get_value(measure(m_poly,U))
22 println("plaq: \$plaq")
23 println("poly: \$poly")
24
25 m_energy = Energy_density_measurement(U)
26 m_topo = Topological_charge_measurement(U)
27 energy = get_value(measure(m_energy,U))
28 topo = get_value(measure(m_topo,U))
29 println("energy: \$energy")
30 println("topo: \$topo")
31
32 m_wilson = Wilson_loop_measurement(U,printvalues=true)
33 wilsonloop = get_value(measure(m_wilson,U))
34 println("wilson loop: ",wilsonloop)
35
36 m_pion = Pion_correlator_measurement(U)
37 m_pion_Staggered = Pion_correlator_measurement(U,fermiontype = "Staggered")
38 m_pion_Wilson = Pion_correlator_measurement(U,fermiontype = "Wilson")
39 pion = get_value(measure(m_pion,U))
40 pion_s = get_value(measure(m_pion_Staggered,U))
41 pion_w = get_value(measure(m_pion_Wilson,U))
42
43 println("pion: \$pion")
44 println("pion correlator with Staggered fermion: \$pion_s")
45 println("pion correlator with Wilson fermion: \$pion_w")
46
47 m_chiral_Staggered = Chiral_condensate_measurement(U,fermiontype = "Staggered")
48 m_chiral_Wilson = Chiral_condensate_measurement(U,fermiontype = "Wilson")
49 chiral_s = get_value(measure(m_chiral_Staggered,U))
50 chiral_w = get_value(measure(m_chiral_Wilson,U))
51
52 println("Chiral condensate with Staggered fermion: \$chiral_s")
53 println("Chiral condensatewith Wilson fermion: \$chiral_w")
54
55
56 TC_methods = ["plaquette","clover"]
57 m_topo = Topological_charge_measurement(U,TC_methods = TC_methods)
58 g = Gradientflow(U)
59 for itrj=1:100
60 flow!(U,g)
61 @time plaq_t = get_value(measure(m_plaq,U))
62 @time poly = get_value(measure(m_poly,U))
63 println("\$itrj plaq_t = \$plaq_t")
64 println("\$itrj polyakov loop = \$(real(poly)) \$(imag(poly))")
65
66 @time topo = get_value(measure(m_topo,U))
67 print("\$itrj topological charge: ")
68 for (key,value) in topo
[…]
70 end
[…]
72 end
73
74end
75test()
One can also use the dictionary type:
1using QCDMeasurements
2using Gaugefields
3function SU3test()
4 println("SU3test")
5 NX = 4
6 NY = 4
7 NZ = 4
8 NT = 4
9 Nwing = 0
10 Dim = 4
11 NC = 3
12
13 U = Initialize_4DGaugefields(NC,Nwing,NX,NY,NZ,NT,condition = "cold")
14 filename = "testconf.txt"
15 L = [NX,NY,NZ,NT]
16 load_BridgeText!(filename,U,L,NC)
17
18 method = Dict()
19 methodname = "Eigenvalue"
20 method["methodname"] = methodname
21 method["fermiontype"] = "Wilson"
22 $\kappa$ = 0.141139
23 method["hop"] = $\kappa$
24 method["nev"] = 1 #number of eigenvalues
25 m = prepare_measurement_from_dict(U,method)
26 value,vectors = get_value(measure(m,U)) #eigenvalues and eigenvectors
27 println("\$methodname \$value")
28
29
30 method = Dict()
31 methodname = "Pion_correlator"
32 method["methodname"] = methodname
33 method["fermiontype"] = "Staggered"
34 method["mass"] = 1
35 method["Nf"] = 4
36 m = prepare_measurement_from_dict(U,method)
37 value = get_value(measure(m,U))
38 println("\$methodname \$value")
39
40 method = Dict()
41 methodname = "Pion_correlator"
42 method["methodname"] = methodname
43 method["fermiontype"] = "Wilson"
44 method["hop"] = 1
45 m = prepare_measurement_from_dict(U,method)
46 value = get_value(measure(m,U))
47 println("\$methodname \$value")
48
49
50 methodsname = ["Plaquette","Polyakov_loop","Topological_charge","Chiral_condensate",
51 "Pion_correlator","Energy_density","Wilson_loop","Eigenvalue"]
52 method = Dict()
53 for methodname in methodsname
54 method["methodname"] = methodname
55 m = prepare_measurement_from_dict(U,method)
56 value = get_value(measure(m,U))
57 if methodname == "Eigenvalue"
58 println("\$methodname \$(value[1])")
59 else
60 println("\$methodname \$(value)")
61 end
62 end
63
64end
65SU3test()
Appendix E Consistency check
We compared results of LatticeQCD.jl to the following papers and codes
  • $N_{f}=4$ SU(3) staggered HMC Gavai et al. (1989)
  • Quenched SU(2) improved thermodynamics Giudice and Piemonte (2017)
  • RHMC Tomiya et al. (2018)
  • HMC for Wilson and Clover Wilson fermions Lattice Tool Kit
  • Pion correlator with the Wilson-Dirac operator Butler et al. (1994)
  • Pion correlator with the staggered Dirac operator Bowler et al. (1988)